OpenStudy (anonymous):

Tension in the rope at the rigid end is (g=10m/s^2) -----|----- | 60kg O (Upward Acceleration 2m/s^2) | 50kg O (Downward 2m/s) | 40kg O (Downward accelration 1m/s^2) OPTIONS: A) 760N B)1360N C)1580N D)1620N

OpenStudy (anonymous):

I MEANT rope

OpenStudy (irishboy123):

if you want help, maybe scan and post the original question or draw something helpful :p acceleration and rope?

OpenStudy (anonymous):

k one sec

OpenStudy (anonymous):

K here is the original question

OpenStudy (anonymous):

FYI the answer is 1580N I just dunnoe how! :((((((((((

OpenStudy (irishboy123):

ok, try it as a simple static mass question. but adjust gravity for each climber so for the first one, accelerating upwards at 2m/s, use g = 9.8 +2 for the others, use g = 9.8-2 and g= 9.8 - 1. where does that take you?

OpenStudy (anonymous):

1450N?

OpenStudy (anonymous):

cuz its 11.8 X 60 + 7.8X 50 + 40X 8.8

OpenStudy (anonymous):

but the question lets us approximate g=10m/s^2 :\

OpenStudy (anonymous):

If i take g=10 and do what u said then I get 1480N

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (michele_laino):

here is my reasoning: on the rope is acting the three weight forces of the corresponding people, so the reaction opposed by the rigid end is: (60+40+50)*10=1500 Newtons. Furthermore, the person of 50 Kg is going upward, so the rigid end is applying a force directly downward, which is 60*2=+120 Newtons, the person of 50 Kg is moving by uniform motion, so he is not applying any force on a rope, apart his weight, which is already computed. Finally the person of 40 Kg is going downward, so the rigid end is applying a force directed upward, which is -1*40=-40 Newton. Then the force applied by the rigid end, of the rope is: 1500+120-40=1580 Newtons

OpenStudy (irishboy123):

** the person of 50 Kg is moving by uniform motion** missed that tots. B is measured in \(\frac{m}{s}\), not \(\frac{m}{s^2}\) that means (10 + 2) X 60 + (10 + 0) X 50 + (10 - 1) X 40 = 1580N

OpenStudy (irishboy123):

thanks @Michele_Laino

OpenStudy (irishboy123):

aplos @sahana2397 i need to learn to read

OpenStudy (michele_laino):

thanks! @IrishBoy123

OpenStudy (michele_laino):

you are good in physics! @IrishBoy123

OpenStudy (irishboy123):

@Michele_Laino and you are better :-))

OpenStudy (irishboy123):

much better

OpenStudy (anonymous):

@Michele_Laino and @IrishBoy123 thanks a ton!!!! :)))))) literally broke my head all day abt this!!! I can finally go to sleep!!! You both rock! P.S @IrishBoy123 i need to learn to read properly aswell :P